Eric is piloting a hot air balloon and drifted 12 miles north,,9 miles west,22 miles south, 3 miles east, and 5 miles south.How many miles does Eric need to turn N,E,S,W to return back to his landing site?
Answer by Fombitz(32388)

Let North be positive y, South negative y.
Let East be positive x, West negative x.
Eric starts at (0,0).
12 miles N: (0,0)+(0,12)=(0,12)
9 miles W:(0,12)+(-9,0)=(-9,12)
22 miles S:(-9,12)+(0,-22)=(-9,-10)
3 miles E:(-9,-10)+(3,0)=(-6,-10)
5 miles S:(-6,-10)+(0,-5)=(-6,-15)
In otder to return to his starting point,
(-6,-15)+(X,Y)=(0,0)
.
.
.
X=6 miles East
Y=15 miles North